# Layer-wise Update Aggregation with Recycling for Communication-Efficient Federated Learning

### v1

@@ 3.3 Theoretical Analysis
LBGM (Low-rank Approximation) $54.87\pm 0.5\%$ $0.65$ $57.13\pm 0.2\%$ $0.87$ $69.83\pm 1.0\%$ $0.71$ $77.96\pm 0.1\%$ $0.23$
FedPAQ (Quantization) $57.42\pm 0.2\%$ $0.50$ $36.15\pm 0.1\%$ $0.50$ $71.54\pm 0.1\%$ $0.25$ $82.72\pm 0.1\%$ $0.25$
FedPara (Re-parameterization) $55.16\pm 0.1\%$ $0.51$ $46.14\pm 0.1\%$ $0.61$ $67.69\pm 0.1\%$ $0.22$ $75.22\pm 0.1\%$ $0.69$
PruneFL (Pruning) $56.76\pm 0.1\%$ $0.51$ $59.40\pm 0.1\%$ $0.69$ $69.42\pm 0.4\%$ $0.19$ $77.25\pm 0.1\%$ $0.22$
FedDropoutAvg (Dropout) $56.54\pm 0.3\%$ $0.50$ $51.25\pm 0.1\%$ $0.60$ $70.61\pm 0.1\%$ $0.25$ $64.94\pm 0.1\%$ $0.50$
FedLUAR (Proposed) $\mathbf{60.15\pm 0.7\%}$ $\mathbf{0.47}$ $\mathbf{59.73\pm 0.6\%}$ $\mathbf{0.61}$ $\mathbf{73.17\pm 0.1\%}$ $\mathbf{0.18}$ $\mathbf{82.80\pm 0.1\%}$ $\mathbf{0.17}$
Table 1: The classification performance comparisons. The communication cost is normalized relative to that of full model averaging. The hyper-parameters are tuned to achieve comparable accuracy to FedAvg while maximizing communication cost reduction.
4 Experiments
Experimental Settings – All experiments are conducted on a GPU cluster which has 2 NVIDIA A6000 GPUs per machine. We use TensorFlow 2.15.0 for training and MPI for model aggregations. All individual experiments were performed at least 3 times, and the average accuracies are reported.
Datasets – We evaluate the performance of our proposed method on representative benchmarks: CIFAR-10 (ResNet20 He et al. (2016)), CIFAR-100 (Wide-ResNet28 Zagoruyko and Komodakis (2016)), FEMNIST (CNN), and AG News (DistillBert Sanh et al. (2019)). When tuning hyper-parameters, we conduct a grid search with a sufficiently small unit size (e.g., 0.1 for learning rate). To generate non-IID datasets, we use label-based Dirichlet distributions with $\alpha=0.1$, which indicates highly non-IID conditions.
$\delta$ CIFAR-10 CIFAR-100
Acc. (%) Comm. Acc. (%) Comm.
0 $61.27\pm 0.7\%$ $1.00$ $59.88\pm 0.8\%$ $1.00$
4 $61.25\pm 0.4\%$ $0.84$ $59.85\pm 0.1\%$ $0.88$
8 $60.92\pm 1.7\%$ $0.68$ $59.93\pm 0.1\%$ $0.76$
12 $60.15\pm 0.7\%$ $0.47$ $59.73\pm 0.6\%$ $0.61$
16 $50.07\pm 1.6\%$ $0.30$ $55.03\pm 0.7\%$ $0.51$
20 N/A N/A $49.60\pm 0.2\%$ $0.36$
Table 2: The CIFAR-10/100 (ResNet20/WideResNet28) classification performance with varying $\delta$ settings. The accuracy remains nearly the same until $\delta$ grows to $12$ for both datasets.
4.1 Comparative Study
We begin by presenting an accuracy comparison among SOTA communication-efficient FL methods below.
  • LBGM Azam et al. (2021)
  • FedPAQ Reisizadeh et al. (2020)
  • FedPara Hyeon-Woo et al. (2021)
  • PruneFL Jiang et al. (2022)
  • FedDropoutAvg Gunesli et al. (2021)
Look-back Gradient Multiplier (LBGM) is a communication-efficient FL method which reuses the model updates in a low-rank space. FedPAQ is a quantization-based FL method that reduces the communication cost in model aggregation by quantizing all individual model parameters. FedPara is a low-rank model decomposition-based FL method which is a representative re-parameterization technique. PruneFL tackles the communication cost in FL by dropping a part of model parameters. FedDropoutAvg is a recently proposed FL method that allows each client to randomly drop the parameters at every round.
Table 1 shows the performance comparison (See Appendix). The total number of clients is 128 and randomly chosen 32 clients participate in every communication round. To make fair comparisons, we find algorithm-specific settings that achieve accuracy reasonably close to the baseline (FedAvg) while minimizing communication costs, and then directly compare the validation accuracy across algorithms.
Overall, FedLUAR achieves accuracy comparable to the baseline while significantly reducing communication costs across all four benchmarks. Particularly for FEMNIST and AG News, our method achieves nearly the same accuracy as FedAvg reducing communication cost to less than $20\%$. Additionally, our proposed method outperforms all the other SOTA methods. Regardless of the dataset, FedLUAR consistently achieves the best accuracy among all the communication-efficient FL methods while having a remarkably reduced communication cost. These results clearly demonstrate that the proposed LUAR method effectively identifies the least critical layers and enables the recycling of previous updates, thereby minimizing communication costs while maintaining model accuracy.
@@ 4.1 Comparative Study
Figure 3: The number of model aggregations per layer. In all the benchmarks, FedLUAR dramatically reduces the number of model aggregations in each layer. The difference between the FedAvg count and the FedLUAR count shows how many times the updates were recycled (how many communications were skipped).
4.2 Harmonization with Other FL Methods
The proposed FL method does not have any dependencies on local training algorithms. To demonstrate this, we apply LUAR to advanced FL algorithms including FedProx Li et al. (2020), MOON Qinbin Li (2021), FedOpt Reddi et al. (2021), FedMut Hu et al. (2024), and FedACG Geeho Kim (2024), and analyze its impact on model accuracy. Table 4 shows FEMNIST accuracy comparisons (See Appendix for the detailed settings). LUAR maintains the validation accuracy while significantly reducing the communication cost across all three benchmarks. Note that the Comm values are communication cost relative to the full model averaging cost. For example, FedPAQ reduces the communication cost to $50\%$ compared to FedAvg, and LUAR further reduces it to $22\%$ of FedPAQ’s communication cost, which results in $11\%$ of FedAvg’s cost. However, LUAR barely affects the accuracy, achieving comparable accuracy to the periodic averaging. These results demonstrate that the proposed method can complement advanced FL methods and be readily applied to many real-world FL applications.
@@ 4.2 Harmonization with Other FL Methods
Table 3: The CIFAR10 performance comparison between before and after applying LUAR to a half layers in ResNet20. LUAR remarkably reduces the communication cost while maintaining the accuracy.
@@
Table 4: The FEMNIST performance comparison between before and after applying LUAR. The communication cost is the ratio of LUAR’s cost to the periodic full averaging cost. The CNN model contains 4 layers in total, and $\delta$ is set to a half of them. For all FL methods, LUAR effectively reduces the communication cost while maintaining the validation accuracy.
@@ 4.3 Communication Cost Analysis
FedLUAR allows clients to skip sending updates for less important layers to the server, thereby reducing communication costs. Figure 3 shows the number of communications for each layer. The more frequently the updates are recycled, the lower the communication cost will be. The communication count charts demonstrate that FedLUAR performs significantly fewer communications than the vanilla FedAvg while achieving nearly the same model accuracy. One intriguing observation is that, in FEMNIST and AG News, the layer with the largest number of parameters is recycled the most, leading to a significant reduction in total communication cost. However, this trend is not observed in the CIFAR-10 and CIFAR-100 benchmarks. Thus, we can conclude that our proposed algorithm is independent of layer size or specific model architectures, adaptively identifying the least significant layers regardless of the model design.
4.4 Additional Performance Analysis
@@
How much does it accelerate? – Figure 4 shows CIFAR-10 learning curves. The x-axis is a relative communication cost to FedAvg. To highlight the difference clearly, we selectively present comparisons among four methods only. The comparison clearly shows that FedLUAR achieves similar accuracy to FedAvg much faster than other SOTA communication-efficient FL methods. Since our method incurs little to no additional computational cost, the same performance gain can be expected in terms of the end-to-end training time in realistic FL environments. See Appendix for more curve charts.
How much could be recycled safely? – We also investigate the impact of $\delta$ on the model accuracy as well as the communication cost. We divide the number of model aggregations at each layer by the total number of communication rounds to calculate the layer-wise communication cost. Then, we sum up the calculated layer-wise costs to get the total communication cost. Intuitively, the larger the $\delta$, the lower the communication cost. However, the model accuracy is expected to drop as more layers recycle their updates.
Table 2 shows CIFAR-10 and CIFAR-100 experimental results for various $\delta$ settings. One key observation is that the accuracy is almost not degraded when LUAR is applied with $\delta\leq 12$ for both datasets. This means that many network layers have quite stable gradient dynamics, and thus their updates can be safely recycled. In addition, the accuracy is hardly reduced until the communication cost is reduced by almost $50\%$. This is a significant benefit especially in FL environments where the network bandwidth is extremely limited. See Appendix for other benchmark results.
Figure 4: The learning curve comparisons for CIFAR-10 (ResNet20). The x-axis represents the communication cost relative to FedAvg. To highlight the differences clearly, we selectively present comparisons among four different methods only.
@@ 5 Conclusion
In this paper, we demonstrated that selectively recycling updates in specific layers can reduce communication costs in FL while preserving model accuracy. In particular, our study empirically proved that the gradient-to-weight magnitude ratio can serve as a practical metric for identifying the least significant layers. This layer-wise partial model aggregation scheme is expected to facilitate the development of efficient FL applications and promote the partial model training paradigm across various deep learning fields. We consider developing a communication-efficient Large Language Model fine-tuning method based on the update recycling scheme as a promising direction for future work.

### v2

@@ 4.1 Comparative Study
LBGM $54.87\pm 0.5\%$ $0.65$ $57.13\pm 0.2\%$ $0.87$ $69.83\pm 1.0\%$ $0.71$ $77.96\pm 0.1\%$ $0.23$
FedPAQ $57.42\pm 0.2\%$ $0.50$ $36.15\pm 0.1\%$ $0.50$ $71.54\pm 0.1\%$ $0.25$ $82.72\pm 0.1\%$ $0.25$
FedPara $55.16\pm 0.1\%$ $0.51$ $46.14\pm 0.1\%$ $0.61$ $67.69\pm 0.1\%$ $0.22$ $75.22\pm 0.1\%$ $0.69$
PruneFL $56.76\pm 0.1\%$ $0.51$ $59.40\pm 0.1\%$ $0.69$ $69.42\pm 0.4\%$ $0.19$ $77.25\pm 0.1\%$ $0.22$
FDA $56.54\pm 0.3\%$ $0.50$ $51.25\pm 0.1\%$ $0.60$ $70.61\pm 0.1\%$ $0.25$ $64.94\pm 0.1\%$ $0.50$
FedBAT $39.56\pm 0.1\%$ $0.03$ $47.24\pm 0.1\%$ $0.03$ $68.27\pm 0.1\%$ $0.03$ $76.38\pm 0.1\%$ $0.57$
[…]
Table 2: Classification performance comparison. Comm denotes the communication cost relative to FedAvg.
@@
Figure 3: Number of model aggregations per layer. FedLUAR significantly reduces aggregation frequency across all benchmarks. The gap from FedAvg indicates how often updates were recycled (i.e., communications skipped).
@@
PruneFL $6.76\pm 0.1\% 55$$5.43\pm 0.1\%$ $0.49$
((a)) CIFAR-10 (ResNet20)
@@
PruneFL $69.42\pm 0.1\%$ $69.11\pm 0.1\%$ $0.11$
((b)) FEMNIST (CNN)
Table 3: CIFAR-10 and FEMNIST performance comparison between before and after applying LUAR. LUAR is applied to half of the model layers for both datasets, using ResNet20 for CIFAR-10 and a CNN for FEMNIST. The Comm column shows the ratio of LUAR’s cost to the FedAvg’s cost.
4.2 Harmonization with Other FL Methods
The proposed FL method does not have any dependencies on the local training algorithm. To demonstrate this, we apply LUAR to several advanced FL methods, including FedProx [24], MOON [29], FedOpt [30], FedMut [11], FedACG [6], and PruneFL [13], and analyze its effect on model accuracy. Table 3 shows CIFAR-10 and FEMNIST accuracy comparisons (See Appendix for details). LUAR maintains the validation accuracy while significantly reducing the communication cost across all three benchmarks. Comm denotes communication cost relative to full model averaging. For instance, FedPAQ reduces communication to $50\%$ of FedAvg, while LUAR further reduces it to $22\%$ of FedPAQ, just $11\%$ of FedAvg, without compromising accuracy. These results show that LUAR can effectively complement advanced FL algorithms and is readily applicable to real-world scenarios.
@@ 4.3 Communication Cost Analysis
FedLUAR enables clients to skip uploading updates for less important layers, thereby reducing communication costs. Figure 3 shows the number of communications for each layer. The communication count charts indicate that FedLUAR requires significantly fewer communications than vanilla FedAvg, while achieving comparable model accuracy. An interesting observation is that, in FEMNIST and AG News, the layer with the largest number of parameters tends to be recycled most frequently, resulting in a substantial reduction in total communication cost. However, this trend is not observed in the CIFAR-10 and CIFAR-100 benchmarks. Thus, we conclude that the proposed method is independent of layer size and specific model architectures, as it adaptively identifies the least significant layers regardless of the model design.
Layer Selection Scheme CIFAR-10 FEMNIST AG News
Acc. (%) Comm. Acc. (%) Comm. Acc. (%) Comm.
Random $53.94\%$ 0.48 $71.10\%$ 0.51 $80.27\%$ 0.23
Top (input-side) $56.03\%$ 0.73 N/A $79.71\%$ 0.21
Bottom (output-side) $45.02\%$ 0.24 $69.54\%$ 0.13 $81.14\%$ 0.45
[…]
Deterministic recycling $48.47\%$ 0.20 $69.08\%$ 0.02 $80.22\%$ 0.15
LUAR(Proposed) $\mathbf{60.15}\%$ 0.47 $\mathbf{73.17}\%$ 0.18 $\mathbf{82.80}\%$ 0.17
Table 4: Performance comparison with different layer selection schemes. For CIFAR-10 and FEMNIST, half the layers were reused; for AG News, 30 layers. Comm. denotes communication cost normalized to FedAvg. Selecting top layers in FEMNIST leads to early-stage divergence.
4.4 Ablation Study
To further validate the effectiveness of the proposed metric shown in (1), we conduct an ablation study as follows. Fixing the number of layers to recycle updates, $\delta$, we measure model accuracy using different layer selection metrics. By comparing these accuracies, we can determine which metric is most effective in identifying the least critical layers in terms of their contribution to the global model training. In particular, we compare the classification performance between the most popular gradient-based layer selection and our proposed LUAR. Table 4 shows the performance comparisons.
This ablation study provides several key insights. First, LUAR outperforms uniform random sampling, demonstrating that our proposed metric (1) effectively identifies less critical layers. Second, even with the same metric, a deterministic selection strategy yields lower accuracy. Persistently recycling updates for layers with low $s_{t,l}$ values can cause them to be too much outdated, introducing excessive noise that degrades model performance. Third, LUAR consistently outperforms the gradient norm-based method, supporting our earlier observation (Fig. 1) that gradient magnitude alone is insufficient to assess update importance. We thus conclude that the gradient-to-weight ratio best captures update quality, achieving the highest accuracy while significantly reducing communication costs.
Dataset Dropping Recycling Comm. Cost $\delta$
CIFAR-10 $46.89\pm 0.1\%$ $50.07\pm 1.6\%$ $0.30$ 16
FEMNIST $64.69\pm 0.2\%$ $73.17\pm 1.1\%$ $0.18$ 2
AG News $77.05\pm 0.1\%$ $82.80\pm 0.1\%$ $0.17$ 30
Table 5: Benchmark performance comparison between update dropping and update recycling schemes.
Additionally, we compare the classification performance of update dropping and recycling schemes. Many existing communication-efficient FL methods merely drop a subset of updates. Table 5 presents the performance comparisons. Here, Dropping refers to the case where the $\delta$ least critical layers are selected using LUAR and their updates are dropped instead of being recycled. As expected, Dropping achieves the same communication cost reduction as Recycling, but its accuracy is significantly lower than that of Recycling. This ablation study clearly demonstrates the superior performance of our proposed update recycling scheme.
((a)) CIFAR-10
((b)) AG News
Figure 4: Learning curves for CIFAR-10 (ResNet20) and AG News (DistillBERT), with communication cost (x-axis) normalized to FedAvg. Four representative methods are shown for clarity.
@@ 4.4 Ablation Study
How much does it accelerate? – Figure 4 shows the learning curves for CIFAR-10 and AG News. The x-axis represents the communication cost relative to FedAvg. To highlight the difference clearly, we selectively present comparisons among four methods only. The comparison clearly shows that FedLUAR achieves similar accuracy to FedAvg much faster than other SOTA communication-efficient FL methods. Since our method incurs little to no additional computational cost, the same performance gain can be expected in terms of the end-to-end training time in realistic FL environments. In our empirical study, we observed the same performance gains across many different FL benchmarks. See Appendix for more curve charts and the detailed experimental settings.
@@ 5 Conclusion
In this paper, we demonstrated that selectively recycling updates in specific layers can reduce communication costs in FL while preserving model accuracy. In particular, our study empirically proved that the gradient-to-weight magnitude ratio can serve as a practical metric for identifying the least significant layers. This layer-wise partial model aggregation scheme is expected to facilitate the development of efficient FL applications and promote the partial model training paradigm across various deep learning fields. We consider developing a communication-efficient Large Language Model fine-tuning method based on the update recycling scheme as a promising direction for future work. A discussion of the broader impact of this work is provided in Appendix A.1.
Acknowledgments
This work was partly supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.RS-2022-00155915, Artificial Intelligence Convergence Innovation Human Resources Development (Inha University)) and National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT)(No. RS-2024-00452914).
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly describe our main contributions, including the proposed layer-wise update recycling method and its communication efficiency benefits, which are consistent with the results shown in Section 4 and 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We give the limitations in Section 3 and Appendix A.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: The theoretical results are fully supported with all assumptions explicitly stated in Section 3.3. Complete and correct proofs are provided in Appendix A.2.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We report all settings and details necessary to reproduce the main experimental results in Section 4 and Appendix A.3.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code is publicly available at https://github.com/swblaster/FedLUAR, and the repository link is also provided at the bottom of page 2 of the paper for reference.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We report all settings and details of experiments in Section 4 and Appendix A.3.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We do not report error bars, as each experiment was repeated 3 times and the mean result is presented.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Section 4 and Appendix A.3 provide sufficient information on the computer resources required to reproduce the experiments.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The research fully adheres to the NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We report broader impacts in Appendix A.1.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: This paper does not involve the release of any pretrained models, data, or systems that carry a high risk of misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We will make the code public upon acceptance of the paper, and the models and datasets used are all publicly available, so no licenses are required.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This paper does not introduce new assets.
  65. Guidelines
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This paper does not involve any crowdsourcing or research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This paper does not involve any human subjects and therefore does not require IRB approval.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: This paper does not use LLMs in any important, original, or non-standard way as part of the core methodology.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.